# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ 3.2 Description of the “moving-blob” model
where $t^{o}_{\mathrm{obs}}$ is the total duration of the observations in the Earth frame. Fitting the model SED to the data, we keep $t^{o}_{\mathrm{shift}}$, $x_{s}$, $x_{f}$ as free parameters, whereas $\theta_{\mathrm{view}}$ is fixed, and $t^{o}_{\mathrm{obs}}$ is also fixed equal to the actual duration of observations over which the observed SED is averaged (7 d if not stated otherwise). Thus, $\beta$, hence $\Gamma$ and $\delta$ as well, become dependent variables following
@@
In the AM3 simulation, we set the number of time steps equal to $N_{s}=100$. At each time step the current position of the blob is updated according to the equation (see Eqs. 27 and 28)
@@
All the external parameters dependent on $x$ are updated at the beginning of each step. The photon fields are calculated the same way it is described in Sect. 3.1.1. The escape timescales are calculated the same way as in Sect. 3.1.2. The snapshot of the model SED$(E;t^{\prime}_{k})$ is stored at each $k$th time step. Then, for each $j$th day of the flare (from first to seventh), following Eqs. (29) and (30), we calculate the model SED averaged over all those times $t^{\prime}_{k}$ in the blob frame which satisfy the condition
@@
where $\Theta$ denotes the Heaviside step function, $t^{\prime}_{e_{0}}$ is the time when the injection starts changing, $\alpha_{e}$ is a free parameter, $x_{e_{0}}$ is the position of the blob when $t^{\prime}=t^{\prime}_{e_{0}}$. We also allow the blob radius to vary over time according to a similar equation
@@
where $t^{\prime}_{b_{0}}$ denotes the time when the blob radius starts changing, $\alpha_{b}$ is a free parameter, $x_{b_{0}}$ is the position of the blob when $t^{\prime}=t^{\prime}_{b_{0}}$. In total, the “moving-blob” model has 12 free parameters: $L^{\prime}_{e}(0)$, $s_{e}$, $E^{\prime{\mathrm{min}}}_{e}$, $E^{\prime{\mathrm{max}}}_{e}$, $B^{\prime}$, $x_{s}$, $x_{f}$, $R^{\prime}_{b}(0)$, $t^{\prime}_{e_{0}}$, $t^{\prime}_{b_{0}}$, $\alpha_{e}$, and $\alpha_{b}$. The transformation of the SED from the blob frame into the observer frame for a single blob propagating along the jet is described by the same Eq. (6) as for the “standing-feature” model satisfying Eq. (7) as shown by Sikora et al. (1997, appendix B).
@@
3.3 Description of the leptohadronic model
maximum contribution of protons to the observed SED allowed by the data assuming the hadronic contribution in the model serves as a subdominant component of the model SED, in addition to the already obtained leptonic fit.
For the best fit of the “standing-feature” leptonic model averaged over MJD 55517–55524, we show the characteristic timescales in the blob frame in Fig. 7. To minimize the number of free parameters, we assume that the protons are co-accelerated with electrons in a process with an acceleration timescale
where $t^{\prime\mathrm{gyro}}_{p}=t^{\prime\mathrm{gyro}}_{e}$ is defined in Eq. (11). We assume for simplicity that $\xi_{\mathrm{acc}}$ is the same for both electrons and protons (see Appendix D for an investigation of other values of $\xi_{\mathrm{acc}}$ for protons). To estimate the value of $\xi_{\mathrm{acc}}$, we look at the interception of the vertical line corresponding to the maximum electron injection energy and the curve of the total electron energy-loss timescale (see Fig. 7) and plot the dependence of the gyroperiod on the energy multiplied by the corresponding value which is found to be $\xi_{\mathrm{acc}}=10^{7}$. This value is the benchmark value for the “standing-feature” leptohadronic model.
To determine the maximum proton energy, we assume that the acceleration started long before the flare [^16] Otherwise, the maximum proton energy would be determined by the interception of $t^{\prime\mathrm{acc}}_{p}$ with the simulation timescale $t^{\prime}_{\mathrm{sim}}$ and would be a few GeV. and look at the interception of the line of $t^{\prime\mathrm{acc}}_{p}(E^{\prime}_{p})$ with the proton escape timescale (and energy-loss timescale due to photopion production having almost the same value at this energy, see Fig. 7):
$t^{\prime\mathrm{acc}}_{p}(E^{\prime{\mathrm{max}}}_{p})=t^{\prime\mathrm{esc}}_{p},$ (38)
where $t^{\prime\mathrm{esc}}_{p}=t^{\prime\mathrm{esc}}_{e}$ is defined in Eq. (13). This results in $E^{\prime{\mathrm{max}}}_{p}\approx 2\times 10^{14}\mathrm{\,eV}$ for MJD 55517–55524, where $E^{\prime{\mathrm{max}}}_{p}$ is taken as the energy of the exponential cutoff of the proton spectrum with the minimal proton energy fixed at $E^{\prime\mathrm{min}}_{p}=2\times 10^{9}$ eV, and their power-law spectral index fixed at $s_{p}=2.0$.
drops to $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). We denote the corresponding value of the proton energy density $u^{\prime\mathrm{UL}}_{p}$ as the upper limit (UL). The obtained ratio between the UL proton energy density and injected electron energy density is
The results obtained with the “standing-feature” leptohadronic model are presented in Sect. 4.3. Since the leptonic model within the “moving-blob” approach resulted in a poor description of the data (see Sect. 4.2), we do not develop its leptohadronic counterpart.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
[…]
  5. 5.
    We obtained a good linear fit in the form of $\Phi^{o}_{\nu}=a\Phi^{o}_{\gamma}$, where $a$ is a constant, and $\Phi^{o}_{\gamma}$ is the integral model photon flux for observable $\gamma$-ray energies $E^{o}_{\gamma}\geq 100$ MeV.
  6. 6.
  7. 7.
  8. 8.
The obtained IceCube neutrino light curve and expected mean neutrino yield are presented in Sect. 4.4.
To estimate the neutrino flux observed by IceCube at energies $E^{o}_{\nu}\geq 100$ TeV, we sum up all signalnesses of all the IceCube alert tracks171717excluding those giving a signal in the IceTop detector from the public IceCat-1 catalogue (Abbasi et al., 2023a). This results in $\approx 12$ “pure” astrophysical muon neutrinos observed by IceCube at energies $E^{o}_{\nu}\geq 100$ TeV.
The estimated contribution of the Fermi-LAT FSRQs to the IceCube neutrino flux is presented in Sect. 4.5.
4 Results
@@ 4.1 Results of the “standing-feature” leptonic model
Figure 9: Best-fit values of the free parameters obtained from the fit of the “standing-feature” model to one-day-long quasi-simultaneous SEDs observed during MJD 55517–55524 along with the reduced chi-squared value (circles). In coloured shades, $1\sigma$ (darker) and $2\sigma$ (lighter) uncertainties are shown, while the MCMC median values are shown as black solid horizontal lines from the MCMC fitting performed for the seven-day-averaged dataset (see Appendix B).
The individual observed SEDs from each of the days of MJD 55517–55524 and the best-fit SED model curves are shown in Fig. 8 along with the fit to the whole-seven-day averaged dataset (Fig. 8h) performed in the same manner as for the one-day-long SEDs with $t^{\prime}_{\mathrm{sim}}$ from Eq. (4) calculated for $t^{o}_{\mathrm{av}}=1$ d to keep this parameter fixed. The best-fit values of the parameters for all seven days are shown in Fig. 9. In coloured shades, we show $1\sigma$ (darker) and $2\sigma$ (lighter) uncertainties derived with emcee (Foreman-Mackey et al., 2013) from the Monte-Carlo Markov Chain (MCMC) fitting for the seven-day-averaged period (see Appendix B) while the MCMC median values are shown as black solid horizontal lines.
From Figs. 8-9 we conclude the following:
@@
    The “standing-feature” pure leptonic model provides a good fit to all one-day-averaged SEDs as well as the seven-day-averaged SED with the exception of MJD 55518–55519 (Fig. 8b): due to the under-fluctuation in the second to last Fermi-LAT bin, the quality of the fit worsens. If one were simply to remove that bin and recalculate the chi-squared (without re-optimization), the reduced chi-squared value would become $\chi^{2}_{\mathrm{red}}=2.16$ and a new optimization could improve this value even further. Thus, we argue that for all seven days of the flare, the “standing-feature” pure leptonic model provides a good description of the observed SEDs.
@@
    The contribution of the direct AD radiation to the optical and UV SED is subdominant.
@@
    The X-ray data are well described as the synchrotron self-Compton (SSC) emission of electrons from the part of the electron spectrum below the cooling break (see Fig. 6) where their cooled spectral index is $p_{e}\approx 2$ resulting in an X-ray spectral index of $\Gamma_{\gamma}\approx 3/2$ (see Eq. 24) which is close to the value obtained from fitting with a functional form (see Table 3).
@@
    The $\gamma$-ray data are well described by IC emission of electrons with energies above the cooling break. The IC process uses the BLR photons as target photons and is in the Thomson regime. The Swift-XRT and Fermi-LAT data are connected with emission from the electron IC on the AD radiation coming from behind the blob in the host galaxy rest frame.
@@
    Among the eight free parameters, it is $s_{e}$ which experiences the most significant change from one day to another during the flare. We argue that this is mostly driven by fluctuations in the highest-energy part of the Fermi-LAT SED due to low photon statistics within one day.
@@
    While the electron injection luminosity does not seem to change significantly within MJD 55517–55524, the Lorentz factor does change and shows an increasing trend throughout the flare. This can be explained as follows. During the last three days of the flare, the Fermi-LAT observes higher values of the $\gamma$-ray SED in the two last energy bins which implies that $x$ must increase in order to avoid $\gamma\gamma$ absorption. Increasing $x$ when it is greater than $R_{\mathrm{BLR}}$ leads to a significant drop in the density of the photon field. To compensate for this drop the Lorentz factor $\Gamma$ should be increased accordingly.
@@
    The blob size stays within the $2\sigma$ MCMC uncertainty of the value obtained for the seven-day-averaged dataset. The inferred variability timescale (see Eq. 14) is approximately from $2$ h to $4.5$ h. These values are close to the values of $2.6\pm 1.0$ h obtained by Meyer et al. (2019).
@@
    For all of the days of the flare the condition $R^{\prime}_{b}\Gamma/x\leq 1$ that the blob cross-section does not exceed the cross-section of the jet in the conical approximation is satisfied.
@@ 4.2 Results of the “moving-blob” leptonic model
Figure 10: “Moving-blob” leptonic model SEDs averaged over one day each from the fit to the seven observed one-day-averaged SEDs for MJD 55517–55524.
Figure 11: The photon and neutrino SED with model curves for the leptohadronic “standing-feature” model with a fit to the observed SED averaged over MJD 55517–5524 with proton acceleration parameter $\xi_{\mathrm{acc}}=10^{7}$. Shaded regions correspond to the spread of proton luminosities $0\leq u^{\prime}_{p}\leq u^{\prime\mathrm{UL}}_{p}$, where $u^{\prime\mathrm{UL}}_{p}$ is the UL corresponding to the Bayes factor with respect to the best-fit leptonic model $BF=10^{-2}$ ($\approx 3\sigma$ Gaussian equivalent). The black solid curve with a downward arrow shows the IceCube one-year sensitivity (see Eq. 42) for the UL of 2.44 muon neutrinos between $E^{o\,\mathrm{min}}_{\nu}=10^{14}$ eV and $E^{o\,\mathrm{max}}_{\nu}=10^{16}$ eV.
@@
Using iMINUIT, we fit the SEDs of the “moving-blob” model to the data from seven one-day-averaged observed SEDs using first the SIMPLEX algorithm and MIGRAD. The SEDs resulting from the fit are shown in Fig. 10. The total quality of the obtained fit for the union of the seven SEDs is rather poor: $\chi^{2}_{\mathrm{red}}\approx 6$. While the model manages to describe the changes in the $X$-ray energy range, it fails to do so in the IR-optical-UV range and in the higher-energy part of the Fermi-LAT spectrum. The latter is mainly due to the movement of the blob far away from the BLR photon field resulting in the decrease of the target photon field energy (see Eq. 17) and density which leads to the under-prediction of the number of the $\gamma$-ray photons at higher energies at later days of the flare. Moreover, to compensate for the decrease of the external photon density, $L^{\prime}_{e}$ is increased accordingly. This would overshoot the X-ray data which are explained by SCC with the flux proportional to $L^{\prime 2}_{e}$. To mitigate overshooting of the X-ray data, the fitter increases the initial blob radius $R^{\prime}_{b}(0)$ up to $6\times 10^{16}$ cm to decrease the synchrotron photon density in the blob frame (decreasing as $R^{\prime-3}_{b}$) which provides a good description of the X-ray data but is in tension with the values expected from the typical flare decay timescale of a few hours (see Sect. 3.1.4).

### v2

@@ 3.2 Description of the “moving-blob” model
where $t^{o}_{\mathrm{obs}}$ is the total duration of the observations in the Earth frame. Fitting the model SED to the data, we keep $t^{o}_{\mathrm{shift}}$, $x_{s}$, $x_{f}$ as free parameters, whereas $\theta_{\mathrm{view}}$ is fixed, and $t^{o}_{\mathrm{obs}}$ is also fixed equal to the actual duration of observations over which the observed SED is averaged (7 d if not stated otherwise). Thus, $\beta$, hence $\Gamma$ and $\delta$ as well, become dependent variables following (see Eqs. 26, 28, and 29)
@@
In the AM3 simulation, we set the number of time steps equal to $N_{s}=100$. At each time step the current position of the blob is updated according to the equation (see Eqs. 26 and 27)
@@
Eq. (31) is valid only in the case of the constant Lorentz factor $\Gamma$; we briefly discuss the case of a varying $\Gamma$ in Sect. 4.2.
[…]
@@
where $\Theta$ denotes the Heaviside step function, $t^{\prime}_{e_{0}}$ is the time when the injection starts changing, $\alpha_{e}$ is a free parameter, $x_{e_{0}}$ is the position of the blob when $t^{\prime}=t^{\prime}_{e_{0}}$. This parametrisation is chosen to allow the electron injection rate to adjust to changing external photon fields, with the increasing rate (in case $\alpha_{e}>0$) to compensate for decreasing photon field density of the BLR photons in the blob frame, while the term $\Theta(t^{\prime}-t^{\prime}_{e_{0}})$ allows to adjust when (and if) the change of the electron injection rate happens. We also allow the blob radius to vary over time according to a similar equation
@@
where $t^{\prime}_{b_{0}}$ denotes the time when the blob radius starts changing, $\alpha_{b}$ is a free parameter, $x_{b_{0}}$ is the position of the blob when $t^{\prime}=t^{\prime}_{b_{0}}$. This parametrisation is chosen to reflect the luminal blob expansion in case $\alpha_{b}=1$ or subluminal expansion in case $\alpha_{b}<1$ (cf., e.g, Zacharias, 2023) while $\Theta(t^{\prime}-t^{\prime}_{b_{0}})$ allows to adjust when (and if) the expansion begins.
In total, the “moving-blob” model has 13 free parameters: $L^{\prime}_{e}(0)$, $s_{e}$, $E^{\prime{\mathrm{min}}}_{e}$, $E^{\prime{\mathrm{max}}}_{e}$, $B^{\prime}$, $x_{s}$, $x_{f}$, $R^{\prime}_{b}(0)$, $t^{\prime}_{e_{0}}$, $t^{\prime}_{b_{0}}$, $t^{o}_{\mathrm{shift}}$, $\alpha_{e}$, and $\alpha_{b}$. The transformation of the SED from the blob frame into the observer frame for a single blob propagating along the jet is described by the same Eq. (4) as for the “standing-feature” model satisfying Eq. (5) as shown by Sikora et al. (1997, appendix B).
@@
4 Results of the leptonic models
@@ 4.1 Results of the “standing-feature” leptonic model
Figure 8: Best-fit values and their $1\sigma$ uncertainties of the free parameters obtained from the fit of the “standing-feature” model to one-day-long quasi-simultaneous SEDs observed during MJD 55517–55524, along with the reduced chi-squared values (dashed-line-connected dots with error bars). For the seven-day-averaged SED, the best-fit values are shown as dark horizontal lines with coloured shades denoting $1\sigma$ (darker) and $2\sigma$ (lighter) uncertainties.
Figure 9: Upper panel: the variability timescale inferred from Eq. (12). Lower panel: the numerical values of the expression $R^{\prime}_{b}\Gamma/x$, which must be $\leq 1$ so that the blob cross-section does not exceed the cross-section of the jet in the conical approximation. The best-fit values and uncertainties (propagated from the uncertainties of the independent parameters shown in Fig. 8) are depicted in the same manner as in Fig. 8.
The individual observed SEDs from each of the days of MJD 55517–55524 and the best-fit SED model curves are shown in Fig. 7 along with the fit to the whole-seven-day averaged dataset (Fig. 7h) performed in the same manner as for the one-day-long SEDs with $t^{\prime}_{\mathrm{sim}}$ from Eq. (2) calculated for $t^{o}_{\mathrm{av}}=1$ d to keep this parameter fixed. Details of the fitting procedure using Monte-Carlo Markov Chain (MCMC) with emcee (Foreman-Mackey et al., 2013) are presented in Appendix F. To demonstrate the uncertainty of the model SED, we draw 300 random MCMC states from the MCMC used for the parameter inference and select those MCMC samples which have their model parameter values lying between the 16th and 84th percentiles of each of their marginal posterior one-dimensional distributions. Then, for each of them, we calculate the model SED and superimpose it in Fig. 7 as a thin solid grey curve. The contribution of each photon field to the observable emission due to the IC process is shown in Fig. 7 with dashed lines with superimposed labels, where SSC stands for “synchrotron self-Compton”, AD EC — “external Compton on the AD radiation”, DT EC — “external Compton on the DT radiation”, BLR EC — “external Compton on the BLR photon field”. The best-fit values with $1\sigma$ uncertainty of the parameters for all seven days are shown in Fig. 8. In coloured shades, we show $1\sigma$ (darker) and $2\sigma$ (lighter) uncertainties from the MCMC fitting for the seven-day-averaged period (see Appendix F) while the MCMC best-fit values for the seven-day-averaged fit are shown as black solid horizontal lines.
From Figs. 7-8 we conclude the following:
@@
    The “standing-feature” pure leptonic model provides a good fit to all one-day-averaged SEDs as well as the seven-day-averaged SED with the exception of MJD 55518–55519 (Fig. 7b): due to the under-fluctuation in the second to last Fermi-LAT bin, the quality of the fit worsens. If one were simply to remove that bin and recalculate the chi-squared (without re-optimisation), the reduced chi-squared value would become $\chi^{2}_{\mathrm{red}}=2.33$ and a new optimisation could improve this value even further. Thus, we argue that for all seven days of the flare, the “standing-feature” pure leptonic model provides a good description of the observed SEDs.
@@
    The contribution of the direct AD+BLR radiation to the optical and UV SED is subdominant.
@@
    The X-ray data are well described as the sum of the electron SSC and EC emission on the DT and AD photon fields (similar to, e.g., the leptonic model of Bottacini et al., 2016).
@@
    The $\gamma$-ray data are well described by IC emission of electrons with energies above the cooling break. The IC process uses the BLR photons as target photons and is in the Thomson regime (with the exception of the highest-energy bins, where some Klein-Nishina cross-section corrections may play a role). The Swift-XRT and Fermi-LAT data are connected with emission from the electron IC on the AD radiation coming from behind the blob in the host galaxy rest frame, and, with a greater contribution, the DT radiation, isotropic in the host galaxy rest frame.
@@
    While the electron injection luminosity does not seem to change significantly within MJD 55517–55524, the Lorentz factor does change and shows an increasing trend throughout the flare. This can be explained as follows. During the last three days of the flare, the Fermi-LAT observes higher values of the $\gamma$-ray SED in the two last energy bins, which implies that $x$ must increase in order to avoid $\gamma\gamma$ absorption. Increasing $x$ when it is greater than $R_{\mathrm{BLR}}$ leads to a significant drop in the density of the photon field. To compensate for this drop, the Lorentz factor $\Gamma$ should be increased accordingly.
@@
    The minimum electron injection energy $E^{\prime{\mathrm{min}}}_{e}$ appears to be well constrained, which is mainly driven by the fact that the peak of the $\gamma$-ray SED observed by Fermi-LAT at a few hundred MeV is due to the break of the cooled electron spectrum happening at $E^{\prime{\mathrm{min}}}_{e}$ as we mentioned in Sect. 3.1.6. Changing $E^{\prime{\mathrm{min}}}_{e}$ significantly would result in a poor description of the high-statistics lower-energy Fermi-LAT $\gamma$-ray SED; besides, it would worsen the fit to the X-ray data as well due to the corresponding shift of the SED peak of the DT EC component.
@@
    For each day, the best-fit value of the blob size stays within the $2\sigma$ MCMC uncertainty of the value obtained for the seven-day-averaged dataset (lighter shade in Fig. 8). The inferred variability timescale (see Eq. 12 and the upper panel of Fig. 9) is approximately from $2$ h to $5$ h. These values are close to the values of $2.6\pm 1.0$ h obtained by Meyer et al. (2019).
@@
    For all of the days of the flare and the seven-day-averaged fit, the condition $R^{\prime}_{b}\Gamma/x\leq 1$ that the blob cross-section does not exceed the cross-section of the jet in the conical approximation [^17] We note that during the fit such a constraint was not imposed. is satisfied for values within $\pm 1\sigma$ around the best-fit values (see the lower panel of Fig. 9).
@@ 4.2 Results of the “moving-blob” leptonic model
Figure 10: Upper panel: “moving-blob” leptonic model SEDs averaged over one day each from the fit to the seven observed one-day-averaged SEDs for MJD 55517–55524. Lower panel: fractional residuals, defined as the differences between the observed and model SEDs divided by the observed SEDs (data points for different dates are slightly shifted in the horizontal direction for better visibility).
Figure 11: The observed (data points with error bars) and “moving-blob” model (solid curves) light curves in the optical (upper panel), X-ray (middle panel), and $\gamma$-ray (lower panel) energy bands. See the legend and the text in Sect. 4.2 for details.
Using iMINUIT, we fit the SEDs of the “moving-blob” model to the data from seven one-day-averaged observed SEDs using first the SIMPLEX algorithm and MIGRAD. The SEDs resulting from the fit are shown in Fig. 10. The total quality of the obtained fit for the union of the seven SEDs is rather poor: $\chi^{2}_{\mathrm{red}}\approx 6$. Due to the significantly larger number of free parameters in the “moving-blob” model than in the “standing-feature” model (13 VS 8) and seven times longer simulation timescale, it is computationally expensive to run an MCMC parameter inference for the former, so, given the poor quality of the fit as well, we limit ourselves only to the iMINUIT minimisation.
Given that the “moving-blob” model is time-dependent, this allows us to plot measured light curves of 3C 454.3 during the flare and compare them to the model ones in Fig. 11. In the upper panel of Fig. 11, the differential time-dependent SED values measured in two optical bands by SMARTS (see Sect. 2.3) are shown as data points with error bars. “Moving-blob” model light curves (solid orange (J band) and green (V band) curves) are in poor agreement with the optical data. In the middle panel of Fig. 11, the integral energy flux measured by Swift-XRT in the energy range 0.510.00.5-10.0 keV is shown as red circles with error bars. The X-ray light curve is taken from the “Open Universe” database (Giommi et al., 2019, 2021) and corrected for the X-ray absorption with the seven-day-averaged value of the hydrogen column density $\kappa$ from Table 2. In this range of the electromagnetic spectrum, the “moving-blob” model (see the solid red curve) clearly shows the best agreement with the data comparing to the other bands, albeit it still fails to describe the sharp drop in the X-ray energy flux after MJD 55521, which might be related to the large size of the blob radius in the model (see below). In the lower panel of Fig. 11, the Fermi-LAT light curve from our analysis (see Sect. 2.1 and Table 1) is shown as one-day-averaged photon flux ($E^{o}_{\gamma}\geq 100$ MeV) data points with error bars. The “moving-blob” model (blue solid curve) reproduces the trend of the observed $\gamma$-ray light curve, although it does not manage to describe the amplitude of the peak of the flare as well as it does for the X-ray data.
@@
While the “moving-blob” model manages to describe the changes in the $X$-ray energy range, it fails to do so in the IR-optical-UV range (significantly over-predicting the optical flux at the earlier times and under-predicting at later times) and in the higher-energy part of the Fermi-LAT spectrum. The latter is mainly due to the movement of the blob far away from the BLR photon field resulting in the decrease of the target photon field energy (see Eq. 15) and density which leads to the under-prediction of the number of the $\gamma$-ray photons at higher energies at later days of the flare (cf. Fig. 7, where, due to nearly constant $x$, the higher-energy part of the Fermi-LAT SED is described by the BLR EC). Moreover, to compensate for the decrease in the external photon density, $L^{\prime}_{e}$ is increased accordingly. This would overshoot the X-ray data which are explained by SCC with the flux proportional to $L^{\prime 2}_{e}$. To mitigate overshooting of the X-ray data, the fitter increases the initial blob radius $R^{\prime}_{b}(0)$ up to $6\times 10^{16}$ cm to decrease the synchrotron photon density in the blob frame (decreasing as $R^{\prime-3}_{b}$) which provides a good description of the X-ray data but is in tension with the values expected from the typical flare decay timescale of a few hours (see Sect. 3.1.4). Moreover, the large size of the blob requires a significant amount of time for electrons to produce synchrotron photons, serving as the target for the SSC process. This results in blob being launched around $t^{o}_{\mathrm{shift}}\approx 2.5$ d earlier than MJD 55517 — the beginning of the period of interest; this causes the over-prediction of the optical flux visible in the upper panel of Fig. 11.
Consideration of a time-dependent (accelerating) blob Lorentz factor $\Gamma$, the hint of which could be seen in Fig. 8 in the “standing-feature” model, does not seem to be able to mitigate the tension between the observed SEDs and the “moving-blob” model SEDs: while in the case of the “standing-feature” the main photon field on which the IC scattering occurs is the deboosted BLR ($x>R_{\mathrm{BLR}}$), the density of which drops drastically with a slight increase of $x$ in later days of the flare, and $\Gamma$ is increased to compensate for this drop, as discussed in Sect. 4.1 (see Fig. 8), the main target for the IC scattering in case of the “moving-blob” model is the DT radiation (the blob propagates from $x_{s}\approx 1.3R_{\mathrm{BLR}}$ to $x_{f}\approx 25R_{\mathrm{BLR}}$), and the blob does not leave the outer radius $R_{\mathrm{DT}}=25R_{\mathrm{BLR}}$ within the timescale of the flare, i.e. the DT photon density remains unchanged if $\Gamma$ is constant. Thus, the accelerating $\Gamma$ would cause shifts of the SEDs to higher energies and an increase of their amplitudes over time, which is not observed. Therefore, we limited ourselves to the case of the constant $\Gamma$ in the “moving-blob” model.